OpenStudy (anonymous):

what is the equation of the line that has a slope of -3 and passes through the point (-6,9)?

hartnn (hartnn):

The equation of line with slope =m and point (x1,y1) is y-y1 = m (x-x1) here, m=-3 and (x1,y1) = (-6,9) just plug in values!